FOURIER TRANSFORM MICROWAVE SPECTROSCOPY OF CuCCH (X^{1}Σ^{+})

M. SUN, \underline{D. T. HALFEN}, Department of Chemistry, Department of Astronomy, and Steward Observatory, University of Arizona, Tucson, AZ 85721; D. J. CLOUTHIER, Department of Chemistry, University of Kentucky, Lexington, KY 40506; and L. M. ZIURYS, Department of Chemistry, Department of Astronomy, and Steward Observatory, University of Arizona, Tucson, AZ 85721.

The rotational spectrum of CuCCH (X^{1}Σ^{+}) has been measured using Fourier transform microwave (FTMW) spectroscopy. This work is the first gas-phase spectroscopic study of this molecule. The species was produced using Discharge Assisted Laser Ablation Spectroscopy (DALAS) in a supersonic jet expansion with HCCH as the precursor molecule. Four rotational transitions (J = 1 → 0, 2 → 1, 3 → 2, and 4 → 3) have been measured for the ^{63}CuCCH and ^{65}CuCCH isotopologues in the range 8-33 GHz. Copper quadrupole splittings have been resolved in both species. The data have been analyzed, and rotational and hyperfine constants determined. Measurements of the carbon-13 and deuterium isotopologues are currently being conducted to establish a precise structure for CuCCH. Compounds of the form CuCCR play an important role in organic synthesis.